OpenStudy (anonymous):

Find the center, vertices, foci, and asymptotes of the hyperbola. x^2/9 - y^2/16=1

OpenStudy (anonymous):

Formula for hyperbola is: \[(x-h)^2/a^2 - (y-k)^2/b^2 =1 \] Center is (0,0) because (h,k) is the center.

OpenStudy (anonymous):

Asymptotes will be + b/a * x and - b/a * x Vertices: (+/- a, 0) Foci: (+/- c, 0) where c^2 = a^2+b^2

OpenStudy (anonymous):